how much pure acid must be added to 6 milliliters of a 5% acid solution to produce a 40% acid solution?
Answer by checkley71(8403)

.05*6+X=.4(6+X) OR .3+X=2.4+.4X OR .6X=2.1 OR X=2.1/6 OR X=3.5 LITERS OF PURE ACID NEEDS TO BE ADDED.